# Some applications of canonical metrics to Landau-Ginzburg models

### v1

@@ 1 Introduction
In the context of Calabi-Yau manifolds, the mirror map, in one direction, sends (complexified) Kähler classes on $X$ to complex structures on the mirror $X^{\vee}$, where both $X$ and $X^{\vee}$ are compact Calabi-Yau. By a classical result of Schumacher [30], there is a separated coarse moduli space of such complex structures on $X^{\vee}$ compatible with any fixed Kähler class, which moreover carries a natural Weil-Petersson metric. It is interesting to study the pullback of this metric to the complexified Kähler cone (see e.g. the discussion by Wilson [39], Section 1).
The situation for Fano manifolds is rather different. The mirror to a Fano $X$ is a Landau-Ginzburg model $Y$, i.e. a noncompact Calabi-Yau manifold endowed with a noncostant holomorphic function $f\!:Y\to\mathbb{C}$. Such objects do not have a good moduli theory in general. As a consequence, compactified Landau-Ginzburg models play an important role in mirror symmetry for Fano manifolds, see for example the discussion by Katzarkov, Kontsevich and Pantev in [24], Section 1, and the related results by Cheltsov and Przyjalkowski [6]. Here we concentrate on the notion of log Calabi-Yau compactification of a toric Landau-Ginzburg model explained e.g. in [29], Definitions 6 and 7.
@@ 1.1 Compactified toric LG models
Fix a pair $(X,H)$ given by a Fano $X$ and an effective divisor $H$. Recall that a log Calabi-Yau compactification of a toric Landau-Ginzburg model (LG model) for $(X,H)$ as above is a projective family $f\!:Z\to\mathbb{P}^{1}$ such that $Z$ is smooth and $-K_{Z}\sim f^{-1}(\infty)$. Moreover, setting $Y:=Z\setminus f^{-1}(\infty)$, there is a dense open torus $(\mathbb{C}^{*})^{n}\subset Y$ on which $f$ restricts to a toric LG model, that is, a Laurent polynomial such that its classical period,
@@
equals the constant term of the regularised $I$-series for $(X,H)$ in genus 00 (a generating function of Gromov-Witten invariants),
$\widetilde{I}^{X,-K_{X},H}_{0}(t):=1+\sum_{\beta,a\geq 0}(-K_{X}\cdot\beta)!\langle\tau_{a}\textbf{1}\rangle_{\beta}e^{-\beta\cdot H}t^{-K_{X}\cdot\beta},$
@@
Finally, the toric LG model should arise from a corresponding toric degeneration of $X$: the Fano manifold $X$ should specialise to a toric variety $T_{X}$ such that the fan polytope $F(T_{X})$ equals the Newton polytope of $f$.
@@
The explicit form of $\widetilde{I}^{X,-K_{X},H}_{0}(t)$ shows that it corresponds to the “polarisation” (big divisor) $-K_{X}+H$. It is standard to extend the above definition to allow a complexified effective divisor $H$ and to regard $-K_{X}+H$ as a generalisation of a complexified Kähler class $[\omega_{\mathbb{C}}]$.
@@
A corresponding expected categorical statement is given by the equivalence
$\operatorname{Fuk}(X,\omega_{\mathbb{C}})\cong\operatorname{MF}(Y,f)\cong\coprod_{\lambda\in\operatorname{Crit}(f)}D^{b}_{\operatorname{sing}}(Y_{\lambda})$
(see [24], Table 1).
@@ 1.2 Maps of moduli spaces
Suppose $X$ is a Fano manifold with a fixed toric degeneration. In the present work we consider the corresponding “framed mirror map” $\Theta$ taking a complexified Kähler class $[\omega_{\mathbb{C}}]\in\operatorname{Ka}_{\mathbb{C}}(X)$ to (a choice of) its compactified toric LG model $Z$ (in cases when this is well defined). We study the problem of realising $\Theta$ as a map taking “(complexified) Kähler moduli” to “complex moduli” in a precise sense, that is, of constructing $\Theta$ as a holomorphic map
$\Theta\!:\operatorname{Ka}_{\mathbb{C}}(X)\supset U\to\mathfrak{M}_{U}$ (1.1)
@@
Our approach relies on the theory of constant scalar curvature Kähler metrics (cscK metrics). Suppose that $Z$ has discrete automorphisms and that we can choose a fixed class $[\eta]\in\operatorname{Ka}(Z)$ such that, as the complex structure on $Z$ varies (holomorphically) with $[\omega_{\mathbb{C}}]\in U$, $[\eta]$ always admits a (necessarily unique, up to automorphisms [11, 5]) cscK representative. Then, the existence of $\Theta$ as a map of moduli spaces, as above, follows from the following classical result.
@@
Suppose $X$ is a smooth del Pezzo surface with a fixed Gorenstein toric del Pezzo degeneration. Then it is possible to construct the corresponding mirror map $\Theta$ at the level of moduli spaces as in (1.1), where $U\subset\operatorname{Ka}_{\mathbb{C}}(X)$ is an explicit, arbitrarily large open neighbourhood of $c_{1}(X)$ and $\mathfrak{M}_{U}$ is a Hausdorff coarse moduli space of cscK (hence $K$-stable) polarised rational elliptic surfaces. In particular, the domain $U$ carries the $(1,1)$-form $\Theta^{*}\Omega_{WP}$, where $\Omega_{WP}$ is the canonical Weil-Petersson metric on $\mathfrak{M}_{U}$.
The proof is carried out in Sections 2-13, as we explain below.
@@
Once we construct the map $\Theta$ as above, we can keep track of the map $f$ as induced by a section of the sheaf over $\mathfrak{M}_{U}$ with fibre over $Z$ given by $H^{0}(Z,-K_{X})$ (indeed, as we recall below, the map to $\mathbb{P}^{1}$ is given by a canonical extension of the corresponding rational pencil).
@@
Let us outline the basic ingredients for the proof of Theorem 1.5. The fixed Gorenstein toric degeneration $T_{X}=T_{\Delta}$ of $X$, where $\Delta$ is a reflexive Fano polytope, provides a toric LG model. There is a corresponding Gorenstein dual toric variety $T_{\nabla}$, such that $\nabla$ is the polar dual of $\Delta$, with an anticanonical rational pencil $f\!:T_{\nabla}\dashrightarrow\mathbb{P}^{1}$. Following [29], Section 3, the canonical log Calabi-Yau compactification $Z$ of $X$ is obtained by constructing a crepant resolution of the (orbifold) singularities of $T_{\nabla}$, as well as of the base locus $\operatorname{Bs}(f)$ of $f$ (one shows that the two sets are disjoint). In the generic case, $\operatorname{Bs}(f)$ is reduced, and we can use a combination of results on Kähler-Einstein (orbifold) metrics [38, 33] and cscK metrics on blowups and resolutions of orbifold singularities (Arezzo-Pacard type theorems such as [1, 2, 37, 3]) in order to construct a family of cscK classes on $Z$. We emphasise that this holds even in cases when $T_{\nabla}$ does not admit an orbifold Kähler-Einstein metric, by choosing the order of metric blowups (i.e. the corresponding Kähler parameters) suitably. This family of cscK polarisations depends on $[\omega_{\mathbb{C}}]\in\operatorname{Ka}_{\mathbb{C}}(X)$, but by analysing the estimates appearing in the proofs of these results, one can see that it is possible to pick a single cscK class $[\eta]$ which works for all $[\omega_{\mathbb{C}}]$ as long as the distances between the base points and between the base locus and the torus-fixed points of $T_{\nabla}$ are uniformly bounded away from zero.
However, this generic result is not enough for our purposes, since there are strata in $\operatorname{Ka}_{\mathbb{C}}(X)$ such that the base locus $\operatorname{Bs}(f)$ is nonreduced. In many cases, these are the most interesting strata, containing for example the anticanonical class $c_{1}(X)$. By construction, the corresponding complex structure on $Z$ is given by an iterated blowup. We show that it is possible to construct cscK classes on such blowups following the strategy above, and that moreover neighbourhoods of such strata can be glued to the generic locus, by using the deformation theory of cscK metrics ([15, 36]). Again, uniform estimates are essential.
@@ 1.4 HMS pairs
One advantage of the approach using canonical Kähler metrics is that, in the case of discrete automorphisms, their deformation theory is unostructed: a small deformation of complex structure of a cscK manifold is still cscK (in the same polarisation for the underlying smooth manifold), see [15, 36]. This lends some flexibility to our construction. We provide an application of this basic fact to mirror pairs in the sense of homological mirror symmetry (HMS).
Suppose $Z$ is a compactified toric LG model endowed with a cscK metric $\omega_{Z}$, as constructed above. Let $\mathfrak{M}$ denote the Fujiki-Schumacher (Hausdorff) moduli space of cscK complex structures compatible with $\omega_{Z}$ (in particular such polarised manifolds are $K$-stable). The pair $(Z,D:=f^{-1}(\infty))$ is a log Calabi-Yau surface. Gross-Hacking-Keel (see [19], Conjecture 0.13) construct a family of del Pezzo surfaces with smooth anticanonical divisors $(\bar{\mathcal{X}},\mathcal{D})\to S=\{|\exp(2\pi\operatorname{i}\int_{D}(B+\operatorname{i}\omega))|<1\,\forall\text{ effective }D\}\subset\operatorname{Pic}(Z)\otimes\mathbb{C}^{*}$ such that, for $s\in S$, the complex structure on $\bar{\mathcal{X}}$ is mirror to the Kähler form on $Z$ in the sense of HMS, that is, we have $D^{b}(\bar{\mathcal{X}}_{s})\cong\operatorname{FS}(Z\setminus D,\operatorname{i}\omega|_{Z\setminus D})$ (this claim was conditional on certain results later proved by Lai-Zhou [26]). Since the cscK condition is scale invariant, we can apply this result to $Z$ up to replacing $[\omega_{Z}]$ with a suitable scaling (to ensure $|s|<1$). Conversely, under the mirror map for pairs $(\bar{\mathcal{X}},\mathcal{D})$, the fixed complex structure on $Z$ corresponds to a Kähler class $[\omega_{\bar{\mathcal{X}}_{s}}]$.
@@ 1.5 Fano threefolds
Proving an analogue of Theorem 1.5 in the case of smooth Fano threefolds is much more challenging of course. Here we confine ourselves to a study of a few special cases. At present, we are only able to obtain a result for small domain $U$, which is moreover conditional on some conjectural expectations.
Firstly, Seyyedali and Székelyhidi [32] prove the analogue of the Arezzo-Pacard criterion for blowing up along submanifolds of codimension greater than $2$, and conjecture that the same result holds for codimension $2$. Secondly, for each smooth Fano threefold $X$, the Fanosearch programme [14] provides several potential Gorenstein toric degenerations, with corresponding toric Landau-Ginzburg models for the divisor $-K_{X}$, such that the period condition is known to hold. Finally, according to [13], Proposition 1.7 and the conjectural expectation of [13], Section 1.4 (partially proved there in Proposition 1.11 and Appendix K), for all such $X$ there exists a well-defined mirror map in a neighbourhood $U$ of $-K_{X}\in\operatorname{Ka}_{\mathbb{C}}(X)$, such that the LG potential corresponding to a class in $U$ is induced by a deformation of the compactified LG model for $-K_{X}$.
@@
Suppose $X$ is a Fano threefold of type $V_{4}$, $V_{6}$, $V_{8}$, $V_{10}$, $V_{12}$, a $(2,2)$ divisor in $\mathbb{P}^{2}\times\mathbb{P}^{2}$, a $(2,2,2)$ branched cover of $(\mathbb{P}^{1})^{3}$ or a $(1,1,1,1)$ divisor in $(\mathbb{P}^{1})^{4}$. Assume the conjectures above. Then, there exist Gorenstein toric degenerations of $X$ such that the corresponding mirror map $\Theta$ can be constructed at the level of moduli spaces as in (1.1), where $U\subset\operatorname{Ka}_{\mathbb{C}}(X)$ is an open neighbourhood of $c_{1}(X)$, and $\mathfrak{M}$ is a Hausdorff coarse moduli space of cscK (hence $K$-stable) polarised manifolds. In particular, $U$ carries the $(1,1)$-form $\Theta^{*}\Omega_{WP}$.
@@
Our observations in Remarks 1.6 (concerning moduli of maps) and 1.8 (on adiabatic stability) also apply to this result (the latter with reference to the higher dimensional results proved in [21], Theorem 1.1). Lemma 1.9 also generalises, as follows.
Lemma 1.13 (Section 14.1).
@@
The assumption that $T_{\nabla}$ is (orbifold) Kähler-Einstein is much too restrictive (this is clear by comparing to the case of del Pezzo surfaces), and we illustrate this in the missing case $X=V_{10}$. In this case $X$ has a Gorenstein toric degeneration such that the dual variety $T_{\nabla}$ is smooth but $\operatorname{Aut}_{0}(T_{\nabla})$ is not reductive (and so a fortiori $T_{\nabla}$ is not even cscK). Nevertheless, $T_{\nabla}$ is isomorphic to a blowup $\operatorname{Bl}_{E}T_{\nabla^{\prime}}$ where $T_{\nabla^{\prime}}$ is Kähler-Einstein, and we can still apply the Arezzo-Pacard Theorem by blowing up $E$ and the components of $\operatorname{Bs}(f)$ in a suitable order.

### v2

@@ 1 Introduction
In the context of mirror symmetry for Calabi-Yau manifolds, it is expected that the mirror map, in one direction, sends (complexified) Kähler classes on $X$ to complex structures on the mirror $X^{\vee}$, where both $X$ and $X^{\vee}$ are compact Calabi-Yau. By a classical result of Schumacher [35], there is a separated coarse moduli space of such complex structures on $X^{\vee}$ compatible with any fixed Kähler class, which moreover carries a natural Weil-Petersson metric. It is interesting to study the pullback of this metric to the complexified Kähler cone (see e.g. the discussion by Wilson [44], Section 1, and Remark 1.8 below).
The situation for Fano manifolds is rather different. The mirror to a Fano $X$ is expected to be a Landau-Ginzburg model $Y$, i.e. a noncompact Calabi-Yau manifold endowed with a noncostant holomorphic function $f\!:Y\to\mathbb{C}$. Such objects do not have a good moduli theory in general. As a consequence, compactified Landau-Ginzburg models play an important role in mirror symmetry for Fano manifolds, see for example the discussion by Katzarkov, Kontsevich and Pantev in [25], Section 2.1, and the related results by Cheltsov and Przyjalkowski [6]. Here we concentrate on the notion of log Calabi-Yau compactification of a toric Landau-Ginzburg model explained e.g. in [33], Definitions 6 and 7.
@@ 1.1 Compactified toric LG models
Fix a pair $(X,H)$ given by a Fano $X$ and an effective divisor $H$. Following Givental [17], the general principle of Hodge theoretic mirror symmetry for Fano manifolds states that genus 00 Gromov-Witten theory on $(X,H)$ should be equivalent to the computation of the classical periods of a regular function $f$ on a noncompact Calabi-Yau manifold $Y$, with respect to a suitable choice of holomorphic volume form $\Omega$. The pair $(f\!:Y\to\mathbb{C},\Omega)$ is known as a Landau-Ginzburg model (LG model).
The notion of toric LG model, recalled below, selects a particular class of genus 00 Gromov-Witten invariants to turn this principle into a definition. As explained in [8], Section B, these are the $1$-pointed, genus 00 Gromov-Witten invariants with descendent
$\langle\beta\psi^{a}\rangle_{0,1,\beta}:=\int_{[X_{0,1,\beta}]^{\operatorname{vir}}}\operatorname{ev}^{*}(\beta)\cup\psi^{a}\in\mathbb{Q},$
where $X_{0,1,\beta}$ denotes the relevant moduli space of stable maps with $1$ marked point with degree $\beta\in H^{*}(X,\mathbb{Z})$, with virtual fundamental class $[X_{0,1,\beta}]^{\operatorname{vir}}$, $\operatorname{ev}$ denotes evaluation at the image of the marked point, and $\psi\in H^{2}(X_{0,1,\beta})$ denotes the first Chern class of the universal cotangent line bundle over $X_{0,1,\beta}$.
Definition 1.1 ([33], Definitions 6 and 7).
A log Calabi-Yau compactification of a toric LG model for $(X,H)$ is a projective morphism $f\!:Z\to\mathbb{P}^{1}$, satisfying the following conditions
  • $(i)$
    log Calabi-Yau condition: $Z$ is smooth and we have $-K_{Z}\sim f^{-1}(\infty)$;
  • $(ii)$
    period condition: setting $Y:=Z\setminus f^{-1}(\infty)$, there is a dense open torus $(\mathbb{C}^{*})^{n}\subset Y$ on which $f$ restricts to a toric LG model for $(X,H)$, that is, a Laurent polynomial such that its classical period,
@@
    equals the constant term of the regularized $I$-series for $(X,H)$ in genus 00 (also known as the regularised quantum period of $(X,H)$, see [8], Section B),
    $\widetilde{I}^{X,-K_{X},H}_{0}(t):=1+\sum_{\beta\in H_{2}(X,\mathbb{Z})}(-K_{X}\cdot\beta)!\langle[\operatorname{pt}]\psi^{-K_{X}\cdot\beta-2}\rangle_{0,1,\beta}e^{-\beta\cdot H}t^{-K_{X}\cdot\beta},$
@@
    Here, the standard holomorphic volume form on $(\mathbb{C}^{*})^{n}$ is given by
    $\Omega=\left(\frac{1}{2\pi\operatorname{i}}\right)^{n}\frac{dx_{1}}{x_{1}}\wedge\cdots\wedge\frac{dx_{n}}{x_{n}},$
    and the integration cycle is
    $\Gamma=\{|x_{1}|=\cdots=|x_{n}|=1\};$
  • $(iii)$
    toric condition: the toric LG model $f$ appearing in $(ii)$ arises from a corresponding toric degeneration of $X$, i.e. the Fano manifold $X$ specialises to a toric variety $T_{X}$ such that the fan polytope $F(T_{X})$ equals the Newton polytope of $f$.
@@
The explicit form of $\widetilde{I}^{X,-K_{X},H}_{0}(t)$ shows that it corresponds to the “polarisation” (big divisor) $-K_{X}+H$. It is standard to extend the above definition to allow a complexified effective divisor $H$ (i.e. a $\mathbb{C}$-divisor with effective real part) and to regard $-K_{X}+H$ as a generalisation of a complexified Kähler class $[\omega_{\mathbb{C}}]$.
@@
Katzarkov, Kontsevich and Pantev ([25], Section 2.1) explain how the homological mirror symmetry approach extends (conjecturally) to Fano manifolds. In particular, if $(f\!:Y\to\mathbb{C},\Omega)$ is mirror to a Fano $X$ with complexified Kähler form $\omega_{\mathbb{C}}$, then there should be an equivalence
$\operatorname{Fuk}(X,\omega_{\mathbb{C}})\cong\operatorname{MF}(Y,f)$
between the Fukaya category of $(X,\omega_{\mathbb{C}})$ (a symplectic invariant) and the category of matrix factorisations of $(f\!:Y\to\mathbb{C},\Omega)$ (a holomorphic invariant), see [25], Table 1. In the other direction, one expects an equivalence
$D^{b}(X)\cong\operatorname{FS}(f\!:Y\to\mathbb{C},(\omega_{Y})_{\mathbb{C}},\Omega)$
between the bounded derived category of $X$ (a holomorphic invariant) and the Fukaya-Seidel category of $(f\!:Y\to\mathbb{C},\Omega)$ with respect to a suitable choice of complexified Kähler form $(\omega_{Y})_{\mathbb{C}}$ (a symplectic invariant).
The work [25] explains how the homological mirror symmetry approach leads to the study of a natural class of log Calabi-Yau compactifications of $(f\!:Y\to\mathbb{C},\Omega)$, called tame compactifications, for which homological mirror symmetry gives interesting predictions: for example, they should have unobstructed deformations, as well as intrinsic Hodge numbers, matching the Hodge numbers of $X$ (the latter prediction was confirmed for threefolds in [6]).
@@ 1.2 Maps of moduli spaces
Suppose $X$ is a Fano manifold with a fixed toric degeneration. In the present work we consider the corresponding “framed mirror map” $\Theta$ taking a complexified Kähler class $[\omega_{\mathbb{C}}]\in\operatorname{Ka}(X)_{\mathbb{C}}$ to (a choice of) its compactified toric LG model $Z$ (in cases when this is well defined). We study the problem of realising $\Theta$ as a map taking “(complexified) Kähler moduli” to “complex moduli” in a precise sense, that is, of constructing $\Theta$ as a holomorphic map
$\Theta\!:\operatorname{Ka}(X)_{\mathbb{C}}\supset U\to\mathfrak{M}_{U}$ (1.1)
@@
Our approach relies on the theory of constant scalar curvature Kähler metrics (cscK metrics). Suppose we can show that there exists a holomorphic family of log Calabi-Yau compactified mirrors parametrised by $U$ such that all members $Z$ have discrete automorphisms and that we can choose a fixed class $[\eta]\in\operatorname{Ka}(Z)$ such that, as the complex structure on $Z$ varies holomorphically with $[\omega_{\mathbb{C}}]\in U$, $[\eta]$ always admits a (necessarily unique, up to automorphisms [12, 5]) cscK representative. Then the existence of $\Theta$ as a map of moduli spaces, as above, follows from the following classical result.
@@
Suppose $X$ is a smooth del Pezzo surface with a fixed Gorenstein toric del Pezzo degeneration. Then it is possible to construct the corresponding mirror map $\Theta$ at the level of moduli spaces as in (1.1), where $U\subset\operatorname{Ka}(X)_{\mathbb{C}}$ is an explicit, relatively compact but arbitrarily large open neighbourhood of $c_{1}(X)$ and $\mathfrak{M}_{U}$ is a Hausdorff coarse moduli space of cscK (hence $K$-stable) polarised rational elliptic surfaces. In particular, the domain $U$ carries the $(1,1)$-form $\Theta^{*}\Omega_{WP}$, where $\Omega_{WP}$ is the canonical Weil-Petersson metric on $\mathfrak{M}_{U}$.
The proof is carried out in Sections 2-13, following the outline we provide in the present Section.
@@
Once we construct the map $\Theta$ as above, we can keep track of the map $f$ as induced by a section of the sheaf over $\mathfrak{M}_{U}$ with fibre over $Z$ given by $H^{0}(Z,-K_{Z})$ (indeed, as we recall below, the map to $\mathbb{P}^{1}$ is given by a canonical extension of the corresponding rational pencil).
@@
Let us outline the basic ingredients for the proof of Theorem 1.6. The fixed Gorenstein toric degeneration $T_{X}=T_{\Delta}$ of $X$, where $\Delta$ is a reflexive Fano polytope, provides a toric LG model. There is a corresponding Gorenstein dual toric variety $T_{\nabla}$, such that $\nabla$ is the polar dual of $\Delta$, with an anticanonical rational pencil $f\!:T_{\nabla}\dashrightarrow\mathbb{P}^{1}$. Following [33], Section 3, the canonical log Calabi-Yau compactification $Z$ of $X$ is obtained by constructing a crepant resolution of the (orbifold) singularities of $T_{\nabla}$, as well as of the base locus $\operatorname{Bs}(f)$ of $f$ (one shows that the two sets are disjoint). Using explicit formulae for the mirror map, one sees that for $[\omega_{\mathbb{C}}]$ lying in a dense open locus, $\operatorname{Bs}(f)$ is reduced and varies holomorphically with $[\omega_{\mathbb{C}}]$. Then we can use a combination of results on Kähler-Einstein (orbifold) metrics [43, 38] and cscK metrics on blowups and resolutions of orbifold singularities (Arezzo-Pacard type theorems such as [1, 2, 42, 3]) in order to construct a family of cscK classes on $Z$. We emphasise that this holds even in cases when $T_{\nabla}$ does not admit an orbifold Kähler-Einstein metric, by choosing the order of metric blowups (i.e. the corresponding Kähler parameters) suitably. This family of cscK polarisations depends on $[\omega_{\mathbb{C}}]\in\operatorname{Ka}(X)_{\mathbb{C}}$, but by analysing the estimates appearing in the proofs of these results, one can see that it is possible to pick a single cscK class $[\eta]$ which works for all $[\omega_{\mathbb{C}}]$ as long as the distances between the base points and between the base locus and the torus-fixed points of $T_{\nabla}$ are uniformly bounded away from zero.
However, this generic result is not enough for our purposes, since there are strata in $\operatorname{Ka}(X)_{\mathbb{C}}$ such that the base locus $\operatorname{Bs}(f)$ is nonreduced. In many cases, these are the most interesting strata, containing for example the anticanonical class $c_{1}(X)$. By construction, the corresponding complex structure on $Z$ is given by an iterated blowup. We show that it is possible to construct cscK classes on such blowups following the strategy above, and that moreover neighbourhoods of such strata can be glued to the generic locus, by using the deformation theory of cscK metrics ([16, 28, 41]). Again, uniform estimates are essential.
@@ 1.4 HMS pairs
One advantage of the approach using canonical Kähler metrics is that, in the case of discrete automorphisms, their deformation theory is unostructed: a small deformation of complex structure of a cscK manifold is still cscK (in the same polarisation for the underlying smooth manifold), see [16, 28, 41]. This lends some flexibility to our construction. We provide an application of this basic fact to mirror pairs in the sense of homological mirror symmetry (HMS).
Suppose $Z$ is a compactified toric LG model endowed with a cscK metric $\omega_{Z}$, as constructed above. Let $\mathfrak{M}$ denote the Fujiki-Schumacher (Hausdorff) moduli space of cscK complex structures compatible with $\omega_{Z}$ (in particular such polarised manifolds are $K$-stable). The pair $(Z,D:=f^{-1}(\infty))$ is a log Calabi-Yau surface. Gross-Hacking-Keel (see [20], Conjecture 0.13) construct a family of del Pezzo surfaces with smooth anticanonical divisors
$(\bar{\mathcal{X}},\mathcal{D})\to S=\{|\exp(2\pi\operatorname{i}\int_{D}(B+\operatorname{i}\omega))|<1,\,\forall\text{ effective }D\}\subset\operatorname{Pic}(Z)\otimes\mathbb{C}^{*}$
($[B+\operatorname{i}\omega]\in\operatorname{Pic}(Z)\otimes\mathbb{C}^{*}$ denoting a point in the complexified Kähler cone) such that, for $s\in S$, the complex structure on $\bar{\mathcal{X}}$ is mirror to the Kähler form on $Z$ in the sense of homological mirror symmetry, that is, we have an equivalence
$D^{b}(\bar{\mathcal{X}}_{s})\cong\operatorname{FS}(Z\setminus D,\operatorname{i}\omega|_{Z\setminus D})$
between the bounded derived category of the category of coherent sheaves on $\bar{\mathcal{X}}_{s}$ and the Fukaya-Seidel category of $(Z\setminus D,\operatorname{i}\omega|_{Z\setminus D})$ (this claim was conditional on certain results later proved by Lai-Zhou [27]). Since the cscK condition is scale invariant, we can apply this result to $Z$ up to replacing $[\omega_{Z}]$ with a suitable scaling (to ensure $|s|<1$). Conversely, under the mirror map for pairs $(\bar{\mathcal{X}},\mathcal{D})$, the fixed complex structure on $Z$ corresponds to a Kähler class $[\omega_{\bar{\mathcal{X}}_{s}}]$.
@@ 1.5 Fano threefolds
Proving an analogue of Theorem 1.6 in the case of smooth Fano threefolds is much more challenging of course. Here we confine ourselves to a study of a few special Fano threefolds $X$. At present, we are only able to obtain a result for a sufficiently small domain $U\ni-K_{X}$, which is moreover conditional on some conjectural expectations.
  • $(a)$
    For each smooth Fano threefold $X$, the Fanosearch programme [15] provides several Laurent polynomials $f_{X}$, such that $f_{X}$ satisfies the period condition (Definition 1.1, $(ii)$) with respect to $-K_{X}$.
    In our applications, given $X$ in our special class, we will make a specific choice of such $f_{X}$ and assume that it actually arises from a toric degeneration of $X$, i.e. that the condition $(iii)$ in Definition 1.1 is satisfied.
  • $(b)$
    Fix $X$, $f_{X}$ as in $(a)$. In our applications, we will fix a log Calabi-Yau compactification $f\!:Z\to\mathbb{P}^{1}$ of $f_{X}$ (constructed as in [33], Section 4).
    We assume that there is a well defined holomorphic map, defined in a neighbourhood $-K_{X}\in U\subset\operatorname{Ka}(X)_{\mathbb{C}}$, which associates to a class $-K_{X}+H\in U$ a log Calabi-Yau compactified LG model for $(X,H)$, obtained as a deformation of $f\!:Z\to\mathbb{P}^{1}$.
    This is the natural modularity condition for a Landau-Ginzburg model, saying that it deforms along with the Kähler parameters of $X$. As we mentioned after Definition 1.1, it is known for del Pezzos, and it is in fact expected for Fano threefolds according to [14], Proposition 1.7 and the conjecture discussed in [14], Section 1.4 (partially proved there in Proposition 1.11 and Appendix K).
  • $(c)$
    Seyyedali and Székelyhidi [37] prove the analogue of the Arezzo-Pacard criterion for blowing up along submanifolds of codimension greater than $2$, and conjecture that the same result holds for codimension $2$.
    In our applications, we assume that this holds, i.e. that [37], Theorem 1 also holds for submanifolds of codimension $2$ of the compactification $Z$ chosen in $(b)$.
@@
Suppose $X$ is a Fano threefold of type $V_{4}$, $V_{6}$, $V_{8}$, $V_{10}$, $V_{12}$, a $(2,2)$ divisor in $\mathbb{P}^{2}\times\mathbb{P}^{2}$, a $(2,2,2)$ branched cover of $(\mathbb{P}^{1})^{3}$ or a $(1,1,1,1)$ divisor in $(\mathbb{P}^{1})^{4}$. Assume that the conjectures $(a)$, $(b)$, $(c)$ above hold for $X$. Then, there exist Gorenstein toric degenerations of $X$ such that the corresponding mirror map $\Theta$ can be constructed at the level of moduli spaces as in (1.1), where $U\subset\operatorname{Ka}(X)_{\mathbb{C}}$ is an open neighbourhood of $c_{1}(X)$, and $\mathfrak{M}$ is a Hausdorff coarse moduli space of cscK (hence $K$-stable) polarised manifolds. In particular, $U$ carries the $(1,1)$-form $\Theta^{*}\Omega_{WP}$.
@@
Our observations in Remarks 1.7 (concerning moduli of maps) and 1.9 (on adiabatic stability) also apply to this result (the latter with reference to the higher dimensional results proved in [22], Theorem 1.1). The Lemma stated in Remark 1.9 also generalises as follows.
Lemma (Section 14.1).
@@
The assumption that $T_{\nabla}$ is (orbifold) Kähler-Einstein is much too restrictive (this is clear by comparing to the case of del Pezzo surfaces), and we illustrate this in the missing case $X=V_{10}$. In this case, $X$ has a Gorenstein toric degeneration such that the dual variety $T_{\nabla}$ is smooth but $\operatorname{Aut}_{0}(T_{\nabla})$ is not reductive (and so a fortiori $T_{\nabla}$ is not even cscK). Nevertheless, $T_{\nabla}$ is isomorphic to a blowup $\operatorname{Bl}_{E}T_{\nabla^{\prime}}$ where $T_{\nabla^{\prime}}$ is Kähler-Einstein, and we can still apply the Arezzo-Pacard Theorem by blowing up $E$ and the components of $\operatorname{Bs}(f)$ in a suitable order.